OpenStudy (anonymous):

WILL GIVE MEDALS! The following table shows the length and width of two rectangles: *Table in comments* Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A? 4x - 10 4x - 14 16x + 18 16x + 14

OpenStudy (anonymous):



OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Well what do we know about the perimeter of a rectangle? P = 2L + 2W If that is true what is the perimeter of rectangle B? P = 2(3x + 1) + 2(2x - 3) P = 6x + 2 + 4x - 6 P = 10x - 4 That is the perimeter of Rectangle B.. So the perimeter of rectangle A would be P = 2(2x + 4) + 2(x - 1) P = 4x + 8 + 2x - 2 P = 6x + 6 The difference of the perimeters would be P = 10x - 4 - (6x + 6) P = 10x - 4 - 6x - 6 P = 4x - 10